# Converting sWeights to Probabilities with Density Ratios

### v1

@@ Abstract
The use of machine learning approaches continues to have many benefits in experimental nuclear and particle physics. One common issue is generating training data which is sufficiently realistic to give reliable results. Here we advocate using real experimental data as the source of training data and demonstrate how one might subtract background contributions through the use of probabilistic weights which can be readily applied to training data. The sPlot formalism is a common tool used to isolate distributions from different sources. However, negative sWeights produced by the sPlot technique can lead to issues in training and poor predictive power. This article demonstrates how density ratio estimation can be applied to convert sWeights to event probabilities, which we call drWeights. The drWeights can then be applied to produce the distributions of interest and are consistent with direct use of the sWeights. This article will also show how decision trees are particular well suited to converting sWeights, with the benefit of fast prediction rates and adaptability to aspects of the experimental data such as data sample size and proportions of different event sources. We also show that a double density ratio approach where the initial drWeights are reweighted by an additional classifier gives substantially better results.
@@ 1 Introduction
The sPlot [1] formalism aims to unfold the contributions of different event sources to the experimental data. The data is assumed to be characterised by discriminating variables for which the distribution of all sources of events are known, and control variables for which the distributions of some or all sources of events are unknown. These control variables are the features that would be used to train a machine learning model.
The sPlot technique is therefore an effective tool to separate various event sources, facilitating the creation of training samples from actual experimental data without the need for detector simulations. However, values of sWeights may be negative [1], which leads to complications when training machine learning algorithms. The weighted binary cross-entropy loss function is defined as:
$L(f(x_{i}))=-\sum\limits_{i}w_{i}(y_{i}\text{log}f(x_{i})-(1-y_{i})\text{log}(1-f(x_{i})))$ (1)
where $f(x_{i})$ is the output of a learning algorithm for event $i$ which has features $x_{i}$, weight $w_{i}$ and label $y_{i}$. In a binary classification task, $y_{i}$ is equal to either zero or one depending on which class the event belongs to, for example $y_{i}$ can be set to one (zero) for a signal (background) event. As learning algorithms are trained to minimise the loss, loss functions must have a lower bound of zero, otherwise the loss can then be made arbitrarily low. For example, given the loss in Equation 1 above, for a negatively weighted signal event where $y_{i}=1$ ($01-y_{i}=$ 0), the output of the learning algorithm could be made arbitrarily close to zero, which would make the loss for that event infinitely small. This is problematic, firstly because the signal event is assigned an output close to zero instead of close to one, and secondly because events with negative weights will then dominate the loss function. In general, loss functions that do not have a lower bound of zero will lead to issues during training and generally poor performance of the learning algorithm.
@@
The key technique here is to recast the sWeights of a given event source to a probability of the event being from that source via the ratio of the density for that source divided by the sum of the densities of all sources. Binary classification can be used to estimate density ratios [6]. Previous works have investigated the use of probability classifier based density ratio estimation. Refs. [7, 8] used a neural network model to calculate weights used for reweighting Monte Carlo samples to more closely agree with data. Ref. [9] used a Boosted Decision Tree model with a bespoke objective function to iteratively reweight histograms. Refs. [10] and [11] used neural networks and boosted decision trees to model the detection efficiency and acceptance of high energy and nuclear physics experiments from simulations of the experiments.
@@
This article will demonstrate how sWeights can be transformed to event probabilities using density ratio estimation whilst preserving the sWeights’ statistical properties. In addition, this article will demonstrate how decision trees are ideally suited to learn sWeights. For certain applications, decision trees are preferable to neural networks as decision trees can have increased computational prediction rates whilst requiring almost no hyperparameter optimisation. Decision trees also typically don’t require as large training datasets as neural networks, which can be beneficial when the amount of experimental data is limited. It is therefore beneficial to have flexibility in the choice of learning algorithm.
The rest of the article is organised as follows: Section 2 will review the sPlot formalism before describing how decision trees cope with negative weights. The remainder of Section 2 will describe how density ratios can be used to convert sWeights to probabilities. Section 3 will then use two case studies, one based on a toy dataset and one based on experimental data taken with the CLAS12 experiment [13], to demonstrate the method’s good performance. Section 4 will end the article with brief conclusions and outlook.
@@ 2.1 Summary of sPlot
The sPlot technique allows to reconstruct variables’ of interest distributions for each event source using the probability density functions (pdf) of each, often established by fitting the expected pdfs on the discriminating variables. In effect, the behavior of the individual sources of events with respect to the variables of interest is inferred from the knowledge available for the discriminating variables by assigning an sWeights to each event in the data sample.
@@ 2.1.1 Decision Trees
One common issue with decision trees is that they are prone to overfitting the training data. One simple solution to reduce overfitting and generally improve the performance of decision trees is so called boosting. The idea behind boosting is that it is generally easier to train several smaller models than a single large model whilst still avoiding over-fitting. Popular boosting algorithms include adaptive boosting [17] or gradient boosting [18]. The boosting algorithm employs multiple decision trees trained one after the other, with subsequent decision trees focused on events incorrectly classified by previous decision trees by assigning a weight to such events. A user defined maximum number of decision trees or a threshold on the training error will stop the boosting process.
The default node splitting loss in the scikit-learn [19] implementation of gradient boosted decision trees is the weighted binary cross-entropy loss. The loss function can be made to avoid issues in training due to negative sWeights by creating a training sample with one class weighted with sWeights and the other with weights set to 1 as defined in Equation 3 and explained in the introduction. The output of a decision tree differs from neural networks in that the output of a leaf depends on the proportion of a class $k$ in a leaf $m$ [20]. For an unweighted classification task, the output is given by:
@@
$\frac{S^{+}_{1}-S^{-}_{1}}{S_{2}}$ $=\frac{S^{+}_{1}}{S_{2}+S^{+}_{2}}$ (6)
$\Rightarrow S^{+}_{2}$ $=\frac{S_{2}S^{-}_{1}}{S^{+}_{1}-S^{-}_{1}}\text{ .}$ (7)
@@ 2.2 Learning Weights using Density Ratios
The aim of the sPlot formalism is to unfold the true distribution of one or more control variables $x_{i}$ for events of different sources using the knowledge available from discriminating variables on the distribution of these sources. The source are classed in species (e.g. signal and background). The sPlot technique provides a consistent representation of how all events from the species are distributed in $x_{i}$ [1]. Summing the sWeights for a given species then recovers the yield of that species obtained by a fit to the discriminating variable. For example, by summing the signal weights one recovers the signal yield. Summing the weights for all species allows to recover the entire data distribution composed of the different species.
@@
$\Rightarrow W_{dr}(x_{i})$ $=\frac{f(x_{i})}{1-f(x_{i})}\text{ .}$ (9)
Overall, the two key aspects of converting sWeights to probabilities using density ratio estimation are first that creating the training sample with one class weighted by the sWeights and the other with weights set to one allows to use the binary cross-entropy loss function without issues in training due to negative sWeights as it preserves a lower bound of zero in the loss function. Second, creating the training sample in such a way allows a binary classification model to learn the ratio of the signal probability density divided by the sum of the probability densities of all species, which is equal to the signal sWeights distribution. Binary classification for density ratio estimation is therefore perfectly suited to convert sWeights to probabilities. Note that the learning algorithm should only be trained with the control variables as inputs, and the learned model works only at the distribution level and not on an event by event basis.
Section 2.1.1 demonstrated how the scitkit-learn implementation of boosted decision trees allows for training with negative weights. In the schema proposed above, negative weights in class 1, with density distribution $D_{S}$, would have the same effect as adding positive weights to class 0, with density distribution $D_{all}$. The density ratio which describes the drWeights $W_{dr}$ is preserved by boosted decision trees, as removing negative weights from $D_{S}$ and adding the corresponding positive weights to $D_{all}$ must preserve the overall ratio.
A final consideration is preserving the uncertainty associated with using the sWeights. This uncertainty is calculated largely by taking the sum of the squared sWeights. However, Ref. [12] showed that calculating the uncertainty as the sum of the resampled weights squared gives incorrect uncertainties. There are two possible solutions, first to just use the sWeights for the uncertainty calculation. Second, Ref. [12] also showed that the uncertainty itself can be converted using density ratio estimation which preserves the uncertainty. For the remainder of this article, we choose the first option, to carry over the uncertainty from the sWeights, but it is important to note that the uncertainty on the sWeights can also be learned.
@@ 3.1 Toy Example
This section will present a toy example to illustrate the performance of the density ratio estimation of sWeights. In the toy example, a simple event generator produces three dimensional events where the first variable is akin to a mass such as the invariant mass of a given reaction, the second variable is akin to an azimuthal ($\phi$) angular distribution and the third variable is $z=\cos{\theta}$. Signal events were generated with a Gaussian distribution in mass and a $\cos{2\phi}$ asymmetry of amplitude 0.8. Background events were generated with a Chebyshev polynomial distribution in mass and a $\cos{2\phi}$ asymmetry of amplitude -0.2. The ratio of signal to background was varied in different tests. The mass variable was used as the discriminatory variable, allowing to separate the signal and background distributions via the weights. The aim of this toy example was then to measure the signal asymmetry in $\phi$ by unfolding the signal and background distributions in the control variables $\phi$ and $z$.
The generated mass distribution was fitted using the sum of the signal and background pdfs used to generate the events with the mean, sigma and polynomial coefficients allowed to vary in the fit. The fit was then used to calculate the signal and background sWeights using the implementation from Ref.[14, 15]. Figure 1 shows a comparison of the mass, $\phi$ and Z distributions for all events and these same events with signal and background sWeights applied, with a signal to background ratio of 1:2. As can be seen, the sum of the signal and background weights allows to reproduced the total distribution. The discriminatory variable, mass, contains negative bins for the signal and background distribution where these events are effectively subtracted to give the disentangled control distributions.
@@
The methodology described in Section 2 was then applied to convert the sWeights using density ratio estimation. The scikit-learn library [19] was used to test both a gradient boosted decision tree (GBDT) and a histogram gradient boosted decision tree (HistGBDT). Both the GBDT and HistGBDT were given a maximum depth of 10 and otherwise default parameters. The neural resampling method described in Ref. [12] was also applied to the toy example. Ref. [12] used particle flow networks (PFN) [21] based on the deep sets architecture [22]. PFNs are general models designed for learning from collider events as unordered, variable-length sets of particles rendering them unsuited to this application with only two variables. Instead neural networks with 7 hidden layers and 1024, 512, 256, 128, 64, 32, 16 nodes respectively were implemented using tensorflow [23]. The hidden layers all had RELU activation functions, with the output layer having a sigmoid activation function, and the network was trained with an ADAM optimiser [24] for 50 epochs. Ref. [11] found that a second iteration of the density ratio estimation, essentially using density ratio estimation for a reweighting step, improved the performance by fine tuning the model. This was also tested here. This second iteration weighted the class labelled 0 with the weights produced by the first density ratio estimation, the predicted weights were then taken as the product of the weights obtained by both individual models.
@@
Training and prediction times were estimated using 5 cores of a AMD EPYC 9554 64-Core Processor at 3.1 GHz. $3.5\times 10^{5}$ events were generated, which leads to a training sample with $7\times 10^{6}$ events as the same events are found in both class 1 and class 0, albeit with and without sWeights weighting. The GBDTs train at a rate of 4kHz and had a prediction rate of roughly 1.6 MHz. The HistGBDTs had a training rate of roughly 1 MHz, with a prediction rate of roughly 20 MHz. The neural network described above had a training rate of 2 kHz with a prediction rate of 0.5 MHz.
@@
The $\phi$ asymmetry was then extracted via the weighted datasets. To do this a 1D histogram was generated in $\phi$ with 100 bins, each bin filled with the relevant signal sWeights or converted weights called drWeights. While the drWeights reproduced the correct distribution, the uncertainties were calculated from the sum of the original sWeights in each bin to allow correct propagation of errors to the fit result as discussed at the end of Section 2.2. Fits were performed by minimizing the $\chi^{2}$ between the model and the binned histogram using the iminuit package [25].
@@
If the sWeights and drWeights accurately separate the signal and background distributions in $\phi$ then the asymmetry obtained by the fit should be consistent with the generated value of 0.8. Several different convertor models were tested along with a double density ratio estimation where the first model was fine tuned by the second model. The total number of generated events was $10^{5}$ with a signal to background ratio of either 1:2 or 1:9. All events were used in training. The entire chain of generating data, fitting and calculating sWeights, training the density ratio model and measuring the $\phi$ asymmetry was repeated 50 times to allow us to determine the robustness of the conversion procedure. Table 1 reports the mean amplitude and uncertainty reported from by the 50 iminuit fits and the standard deviation of the amplitude over the 50 datasets. The expectation is that the mean should be consistent with the nominal value of 0.8, while the mean uncertainty and standard deviation should be numerically similar i.e. the fluctuation of results is consistent with the calculated uncertainty. The first row shows the results for the sWeighted distributions that we are trying to emulate.
Mean $\sigma$ Uncertainty $\frac{\sigma}{\text{Uncertainty}}$ Mean $\sigma$ Uncertainty $\frac{\sigma}{\text{Uncertainty}}$
Signal:Background 1:2 1:2 1:2 1:2 1:9 1:9 1:9 1:9
sWeights 0.802 0.0082 0.0089 0.92 0.804 0.0244 0.0274 0.89
GBDT 0.796 0.0094 0.0091 1.03 0.707 0.0779 0.0271 2.87
GBDT & GBDT 0.807 0.0093 0.0092 1.01 0.793 0.0260 0.0285 0.91
GBDT & HistGBDT 0.810 0.011 0.0092 1.20 0.791 0.0347 0.0283 1.23
HistGBDT 0.760 0.0169 0.0089 1.90 0.643 0.0356 0.0261 1.36
HistGBDT & GBDT 0.788 0.0115 0.0091 1.26 0.739 0.0326 0.0281 1.16
HistGBDT & HistGBDT 0.782 0.0112 0.0090 1.24 0.713 0.0342 0.0274 1.25
NN 0.782 0.0262 0.0089 2.94 0.743 0.0417 0.0264 1.58
NN & GBDT 0.822 0.0130 0.0093 1.40 0.826 0.0369 0.0299 1.23
NN & HistGBDT 0.813 0.0180 0.0093 1.94 0.816 0.0351 0.0291 1.21
Table 1: Comparison of the $\phi$ asymmetry measured for various models used by the density ratio estimation. The signal distribution was generated with a $\phi$ asymmetry of 0.8 for $10^{5}$ events with a signal to background ratio of 1:2 or a signal to background ratio of 1:9. The data generation and training were repeated 50 times. The mean and standard deviation ($\sigma$) of the measured asymmetries are reported, along with the mean uncertainty.
We observe from the results in Table 1 that all convertors except the HistGBDT perform adequately in reproducing the amplitude (mean value) for the 1:2 signal to background case. In particular models with a double density ratio, or reweighting step, are an improvement on the single step case. For the 1:9 case where the background dominates, models with the GBDT reweighting step perform well, while the others do not give as accurate an amplitude. In all cases the uncertainty is consistent as it is just a property of the data statistics and we are using the sWeighted sum of the weights squared. The standard deviation however does vary significantly from the given uncertainty showing a lack of robustness in the procedure for some convertor models. This is likely due to some random events getting inaccurate weights and thereby less accurate $\phi$ distributions. This also seems to be more of an issue for the neural network based models and in the higher background tests. On the other hand the double GBDT model seems to perform admirably even with 1:9 signal to background and would seem to be the most robust convertor.
@@
Figure 2: The total (black) distribution in $\phi$ (left) and Z (right) compared to the same distribution with signal sWeights applied (blue) or the drWeights applied (red). The top row was generated with a signal to background ratio of 1:2, wereas the bottom row has a signal to background ratio of 1:9. The pull between the sWeights and drWeights distributions is also shown below all distributions. The drWeights are estimated using the GBDT and GBDT double density ratio.

### v2

@@ Abstract
The use of machine learning approaches continues to have many benefits in experimental nuclear and particle physics. One common issue is generating training data which is sufficiently realistic to give reliable results. Here we advocate using real experimental data as the source of training data and demonstrate how one might subtract background contributions through the use of probabilistic weights which can be readily applied to training data. The sPlot formalism is a common tool used to isolate distributions from different sources. However, the negative sWeights produced by the sPlot technique can cause training problems and poor predictive power. This article demonstrates how density ratio estimation can be applied to convert sWeights to event probabilities, which we call drWeights. The drWeights can then be applied to produce the distributions of interest and are consistent with direct use of the sWeights. This article will also show how decision trees are particularly well suited to convert sWeights, with the benefit of fast prediction rates and adaptability to aspects of experimental data such as the data sample size and proportions of different event sources. We also show that a density ratio product approach in which the initial drWeights are reweighted by an additional converter gives substantially better results.
[…]
@@ 1 Introduction
The sPlot [1] formalism aims to separate the contributions of different event sources to the experimental data. The data is assumed to be characterized by discriminating variables for which the distribution of all sources of events are known, and control variables for which the distributions of some or all sources of events are unknown. These control variables are the features that would be used to train a machine learning model.
The sPlot technique is therefore an effective tool for separating various event sources, facilitating the creation of training samples from actual experimental data without the need for detector simulations. However, values of sWeights may be negative [1], which leads to complications when training machine learning algorithms. The weighted binary cross-entropy loss function is defined as:
$L(f(x_{i}))=-\sum\limits_{i}w_{i}(y_{i}\text{log}f(x_{i})+(1-y_{i})\text{log}(1-f(x_{i})))$ (1)
where $f(x_{i})$ is the output of a learning algorithm for event $i$ which has features $x_{i}$, weight $w_{i}$ and label $y_{i}$. In a binary classification task, $y_{i}$ is equal to either zero or one depending on which class the event belongs to, for example $y_{i}$ can be set to one (zero) for a signal (background) event. As learning algorithms are trained to minimise the loss, loss functions must have a lower bound, otherwise the loss can then be made arbitrarily low. For example, given the loss in Equation 1 above, for a negatively weighted signal event where $y_{i}=1$ ($1-y_{i}=0$), the output of the learning algorithm could be made arbitrarily close to zero, which would make the loss for that event infinitely small. This is problematic, firstly because the signal event is assigned an output close to zero instead of close to one, and secondly because events with negative weights will then dominate the loss function. In general, loss functions that do not have a lower bound will lead to issues during training and poor performance of the learning algorithm.
@@
The key technique utilised here is to recast the sWeights of a given event source to a probability of the event being from that source via the ratio of the density for that source divided by the sum of the densities of all sources. Binary classification can be used to estimate density ratios [6]. Previous works have investigated the use of probability classifier based density ratio estimation. Refs. [7, 8] used a neural network model to calculate weights used for reweighting Monte Carlo samples to more closely agree with data. Ref. [9] used a Boosted Decision Tree model with a bespoke objective function to iteratively reweight histograms. Refs. [10] and [11] used neural networks and boosted decision trees to model the detection efficiency and acceptance of high energy and nuclear physics experiments from simulations of the experiments.
@@
This article will demonstrate how sWeights can be transformed to event probabilities using density ratio estimation. This article will also demonstrate how decision trees are ideally suited to learn sWeights. For certain applications, decision trees are preferable to neural networks as decision trees can have increased computational prediction rates whilst requiring almost no hyperparameter optimisation. Decision trees also typically do not require as large training datasets as neural networks, which can be beneficial when the amount of experimental data is limited. It is therefore beneficial to have flexibility in the choice of learning algorithm. In addition, this article will discuss the biases and correlations that are introduced in the data when using a learning algorithm to convert negative weights to positive definite probabilities, along with demonstrating that these biases are acceptable for the purposes of creating machine learning training data sets from experimental data.
The rest of the article is organised as follows: Section 2 will review the sPlot formalism before describing how decision trees cope with negative weights. The remainder of Section 2 will describe how density ratios can be used to convert sWeights to probabilities. Section 3 will then use three case studies, two based on a toy data set and one based on experimental data taken with the CLAS12 experiment [13], to demonstrate the good performance of the method. Section 4 ends the article with brief conclusions and perspectives.
@@ 2.1 Summary of sPlot
The sPlot technique allows one to reconstruct variables’ of interest distributions for each event source using the probability density functions (pdf) of each, often established by fitting the expected pdfs on the discriminating variables. In effect, the behavior of the individual sources of events with respect to the variables of interest is inferred from the knowledge available for the discriminating variables by assigning an sWeight to each event in the data sample.
@@ 2.1.1 Decision Trees
One common issue with decision trees is that they are prone to overfitting the training data. One simple solution to reduce overfitting and generally improve the performance of decision trees is so called boosting. The idea behind boosting is that it is generally easier to train several smaller models than a single large model whilst still avoiding over-fitting. Popular boosting algorithms include adaptive boosting [17] or gradient boosting [18]. The boosting algorithm employs multiple decision trees trained one after the other, with subsequent decision trees focused on events incorrectly classified by previous decision trees by assigning a suitably defined weight, $w>1$, to such events. A user defined maximum number of decision trees or a threshold on the training error will stop the boosting process.
The default node splitting loss in the scikit-learn (version 1.5.0) [19] implementation of gradient boosted decision trees is the weighted binary cross-entropy loss. The loss function can be made to avoid issues in training due to negative sWeights by creating a training sample with one class weighted with sWeights and the other with weights set to 1 as defined in Equation 3 and explained in the introduction. The output of a decision tree differs from neural networks in that the output of a leaf depends on the proportion of a class $k$ in a leaf $m$ [20]. For an unweighted classification task, the output is given by:
@@
$\frac{S^{+}_{1}-S^{-}_{1}}{S_{2}}$ $=\frac{S^{+}_{1}}{S_{2}+S^{+}_{2}}$
[…]
@@ 2.2 Learning Weights using Density Ratios
The aim of the sPlot formalism is to separate the true distribution of one or more control variables $x_{i}$ for events of different sources using the knowledge available from discriminating variables on the distribution of these sources. The source are classed in species (e.g. signal and background). The sPlot technique provides a consistent representation of how all events from the species are distributed in $x_{i}$ [1]. Summing the sWeights for a given species then recovers the yield of that species obtained by a fit to the discriminating variable. For example, by summing the signal weights one recovers the signal yield. Summing the weights for all species allows to recover the entire data distribution composed of the different species.
@@
[…]
Overall, the two key aspects of converting sWeights to probabilities using density ratio estimation are first that creating the training sample with one class weighted by the sWeights and the other with weights set to one allows to use the binary cross-entropy loss function without issues in training due to negative sWeights as it preserves a lower bound in the loss function. Second, creating the training sample in such a way allows a binary classification model to learn the ratio of the signal probability density divided by the sum of the probability densities of all species, which is equal to the signal sWeights distribution. Binary classification for density ratio estimation is therefore perfectly suited to convert sWeights into probabilities. Note that the learning algorithm should only be trained with control variables as inputs and not discriminatory variables.
Section 2.1.1 demonstrated how the scitkit-learn implementation of boosted decision trees allows training with negative weights. In the schema proposed above, negative weights in class 1, with density distribution $D_{S}$, would have the same effect as adding positive weights to class 0, with density distribution $D_{all}$. The density ratio which describes drWeights $W_{dr}$ is preserved by boosted decision trees, as removing negative weights from $D_{S}$ and adding the corresponding positive weights to $D_{all}$ must preserve the overall ratio.
A final consideration is to preserve the uncertainty associated with using the sWeights. This uncertainty generally requires taking the sum of the squared sWeights. However, Ref. [12] showed that calculating the uncertainty as the sum of the resampled weights squared gives incorrect uncertainties. There are two possible solutions; the first is to to estimate the uncertainties by the quadratic sum of the sWeights. Second, Ref. [12] also showed that the uncertainty itself can be converted using density ratio estimation, which preserves the uncertainty. For the remainder of this article, we choose the first option, to carry over the uncertainty from sWeights, but it is important to note that the uncertainty on sWeights can also be learned.
@@ 3.1 Toy Example
This section will present a toy example to illustrate the performance of the density ratio estimation of sWeights. In the toy example, a simple event generator produces three dimensional events where the first variable is akin to a mass such as the invariant mass of a given reaction, the second variable is akin to an azimuthal ($\phi$) angular distribution and the third variable is $z=\cos{\theta}$. Signal events were generated with a Gaussian distribution in mass and a $\cos{2\phi}$ of amplitude 0.8. Background events were generated with a linear mass distribution and a $\cos{2\phi}$ of amplitude -0.2. The ratio of signal to background was varied in different tests. The mass variable was used as the discriminatory variable, allowing us to separate the signal and background distributions via the weights. The purpose of this toy example was then to measure the amplitude of the signal in $\phi$ by separating the signal and background distributions in the control variables $\phi$ and $z$.
The generated mass distribution was fitted using the sum of the signal and background pdfs used to generate the events with the mean, width, and polynomial coefficients allowed to vary in the fit. The fit was then used to calculate the signal and background sWeights using the implementation from Ref.[14, 15]. Figure 1 shows a comparison of the mass, $\phi$ and Z distributions for all events and these same events with applied signal and background sWeights, with a signal-to-background ratio of 1:2. As can be seen, the sum of the signal and background weights allows to reproduce the total distribution. The discriminatory variable, mass, contains negative bins for the signal and background distribution, where these events are effectively subtracted to give the disentangled control distributions.
@@
The methodology described in Section 2 was then applied to convert the sWeights using density ratio estimation. The scikit-learn library (version 1.5.0) [19] was used to test both a gradient boosted decision tree (GBDT) and a histogram gradient boosted decision tree (HistGBDT). Both the GBDT and HistGBDT were given a maximum depth of 10 and otherwise default parameters. We found that performance improved with depth up to 10, but not significantly after that. The neural resampling method described in Ref. [12] was also applied to the toy example. Ref. [12] used particle flow networks (PFN) [21] based on the deep sets architecture [22]. PFNs are general models designed for learning from collider events as unordered, variable-length sets of particles, rendering them unsuited to this application with only two variables. Instead, we used neural networks (NN) with 7 hidden layers and 1024, 512, 256, 128, 64, 32, 16 nodes, respectively, implemented using tensorflow (version 2.16.1) [23]. The hidden layers all had RELU activation functions, the output layer had a sigmoid activation function, and the network was trained with an ADAM optimizer [24] for 50 epochs. Ref. [11] found that a second iteration of the density ratio estimation, essentially using the density ratio estimation for a reweighting step, improved the performance by fine-tuning the model. This was also tested here.
The idea behind using a second reweighting step is that after the initial drWeights application there may still be some residual differences between the drWeighted and sWeighted distributions. A second classification may then be applied between the drWeighted and sWeighted distributions. If the two distributions agree, then the weights from this second classification should all be close to 1, but if there are regions where the agreement is not so close, the second weights will correct the first. This second iteration weighted the class labeled 0 with the weights produced by the first density ratio estimation, and the predicted weights were then taken as the product of the weights obtained by both individual models.
@@
Training and prediction times were estimated using 5 cores of an AMD EPYC 9554 64-Core Processor at 3.1 GHz. $3.5\times 10^{5}$ events were generated, leading to a training sample with $7\times 10^{5}$ events since the same events are found in both classes 1 and 0, albeit with and without sWeights weighting. The GBDTs train at a rate of 4kHz and had a prediction rate of roughly 1.6 MHz. The HistGBDTs had a training rate of roughly 1 MHz, with a prediction rate of roughly 20 MHz. The neural network described above had a training rate of 2 kHz with a prediction rate of 0.5 MHz. The impact of the number of training events is discussed later in the text. Unless otherwise mentioned, tests were made with $10^{5}$ generated events, leading to training samples with $2\times 10^{5}$ events since the same events are found in both classes 1 and 0 with or without sWeight weighting.
@@
The $cos(2\phi)$ amplitude was then extracted via the weighted datasets. To do this a 1D histogram was generated in $\phi$ with 100 bins, each bin filled with the relevant signal sWeights or converted weights called drWeights. While the drWeights reproduced the correct distribution, the uncertainties were calculated from the sum of the original sWeights in each bin to allow correct propagation of errors to the fit result as discussed at the end of Section 2.2. Fits were performed by minimizing the $\chi^{2}$ between the model and the binned histogram using the iminuit package [25].
@@
If the sWeights and drWeights accurately separate the signal and background distributions in $\phi$ then the amplitude obtained by the fit should be consistent with the generated value of 0.8. Several different converter models were tested along with a density ratio product estimation where the first model was fine tuned by the second model. The total number of generated events was $10^{5}$ with a signal to background ratio of either 1:2 or 1:9. All events were used in training. The entire chain of generating data, fitting and calculating sWeights, training the density ratio model and measuring the $\cos(2\phi)$ amplitude was repeated 50 times to allow us to determine the robustness of the conversion procedure. We report the mean amplitude and uncertainty $\bar{\sigma}_{fit}$ from the 50 iminuit fits and the standard deviation, $\hat{\sigma}_{rms}$, of the amplitude over the 50 data sets. The expectation is that the mean should be consistent with the nominal value of 0.8, while $\bar{\sigma}_{fit}$ and $\hat{\sigma}_{rms}$ should be numerically similar, i.e. the fluctuation of results is consistent with the calculated uncertainty.
The mean amplitude, $\bar{\sigma}_{fit}$, and $\hat{\sigma}_{rms}$ of the amplitude over 50 iterations when varying several parameters of the drWeights are contained in A for conciseness. Table 3 shows the performance of a single GBDT converter when varying the maximum depth of the GBDT, demonstrating that deeper GBDTs produce better results as they fully capture correlations in the input variable space. A maximum depth of 10 is chosen as further increasing the depth only leads to longer training and prediction times without increasing the performance of the method. Table 4 shows the results from different classifiers such as a HistGBDT, a GBDT and a neural network. The GBDT produces the most reliable results, and in particular adding a reweighting step or density ratio product produces the best performance. Most converters are able to reproduce the nominal amplitude of 0.8 but $\hat{\sigma}_{rms}$ varies significantly from $\bar{\sigma}_{fit}$ for some converter models. However, the density ratio product using two GBDTs is able to reproduce the nominal amplitude of 0.8 for both signal to background ratios of 1:2 and 1:9 and obtains consistent $\hat{\sigma}_{rms}$ and $\bar{\sigma}_{fit}$, proving to be the most robust converter. Table 5 shows the performance of the GBDT density ratio product when varying the number of generated events. In all cases the amplitude measured with the drWeights is consistent with the generated amplitude of 0.8. The $\bar{\sigma}_{fit}$ and $\hat{\sigma}_{rms}$ on the amplitudes are generally consistent. The good performance of the drWeights even with large backgrounds and small event samples is a key consideration, given that experimental datasets may be limited and may have irreducible backgrounds. Although it is worth emphasizing that drWeights should only be used when positive definite probabilities are required as the sWeights proved to be generally more reliable, overall the drWeights are found to be a robust conversion of sWeights to probabilities.
@@
Figure 2: The total (black) distribution in $\phi$ (left) and Z (right) compared to the same distribution with signal sWeights applied (blue) or the drWeights applied (red), for $5\times 10^{5}$ generated events with 0.35M used in training and 0.15M plotted here. The top row was generated with a signal to background ratio of 1:2, whereas the bottom row has a signal to background ratio of 1:9. The pull, $g_{hist}$, between the sWeights and drWeights distributions is also shown below all distributions. The drWeights are estimated using the GBDT and GBDT density ratio product.